OpenStudy (anonymous):

Check my answer? Identifying reflections? Identify a reflection about the x-axis or a reflection about the y-axis of ƒ(x) = x^2 + 4x^3 by observing the equation of the function g(x) = x^2 – 4x^3. (A) A reflection about the x-axis (B) A rreflection about the y-axis I think the answer is B, a reflection about the y axis, can anyone confirm?

OpenStudy (perl):

f(-x) is a reflection of f(x) about the y axis -f(x) is a reflection of f about the x axis

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so my answer is right then @perl ?

ganeshie8 (ganeshie8):

Also explain why you think B is the answer

OpenStudy (anonymous):

4x^3 seems like it would result in a lower number than x^2, and the lower the number means it will travel left across the coordinate grid across the y axis

ganeshie8 (ganeshie8):

you seem to be having right intuition, but this can be worked more efficiently. look at perl's reply

OpenStudy (anonymous):

I'll remember perl's method but is my answer correct?

ganeshie8 (ganeshie8):

f(x) = x^2 + 4x^3 f(-x) = ?

OpenStudy (anonymous):

x^2 - 4x^3

ganeshie8 (ganeshie8):

f(-x) = x^2-4x^3 = g(x) so yeah g(x) is the reflection of f(x) about the y-axis

OpenStudy (anonymous):

thank you

ganeshie8 (ganeshie8):

you're welcome!

OpenStudy (anonymous):

b